OpenStudy (anonymous):

Please help, Algebra 2!

OpenStudy (anonymous):

Where is algebra 2??

OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x – y = 3 7x – 9y = –2 (1, 1) (6, –3) (6, 1) (1, –3)

OpenStudy (anonymous):

the choicees are in parenthesis

OpenStudy (anonymous):

Can you find y from first equation??

OpenStudy (anonymous):

it could be 4(2) - 5= 3 but how do i know what x is?

OpenStudy (anonymous):

Wait it is not the right way to that..

OpenStudy (anonymous):

i dont know :/

OpenStudy (anonymous):

See: 4x - y = 3 So Subtract 4x both the sides: -y = 3 - 4x So now multiply by (-1): y = -3 + 4x or y = 4x - 3

OpenStudy (anonymous):

so x is -1?

OpenStudy (anonymous):

Sorry x = 1 is right..

OpenStudy (anonymous):

so is it (1,-3) ?

OpenStudy (anonymous):

So now you got : y = 4x - 3 Put this value in the second equation in place of y: 7x – 9y = –2 7x - 9(4x-3) = -2 7x - 36x + 27 = -2 -29x = -29 x = 1

OpenStudy (anonymous):

What is : 4 - 3 = ??

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so this is y..

OpenStudy (anonymous):

Which option is this??

OpenStudy (anonymous):

so its 1,1

OpenStudy (anonymous):

Absolutely..

OpenStudy (anonymous):

thankyou very much :)